If the Fathers age is 5 times his son's age. 4 years back the father was 9 times older than son. what is the fathers present age.?
Found 3 solutions by Theo, ikleyn, kiks:
Answer by Theo(13342)

f = father's age
x = his son's age.
f = 5x
f-4 = 9*(x-4)
in the second equation replace f with 5x to get:
5x-4 = 9*(x-4)
simplify to get:
5x-4 = 9x-36
subtract 5x from both sides of the equation and add 36 to both sides of the equation to get:
32 = 4x
solve for x to get:
x = 8
f = 5x means f = 40
f-4 = 36
x-4 = 4
36 = 9 * 4.
solution is confirmed to be good.
solution is father's present age is 40.

Answer by ikleyn(52787)

.

In formal English and in Math problems written in English, such combinations of words

like "9 times older than" and other similar combinations are NOT USED and are PROHIBITED to use,

because they are AMBIGUOUS and provoke misunderstanding.


When a Math composer uses such a combinations of words, it is a 100% sign
that this person does not know (is unfamiliar with) the rules of writing Math problems in English.


It is a 100% sign that this person did not open and did not read a standard Math textbook in English.



Answer by kiks(2)

Let x be the son's age.
s= x
f= 5x
4 years ago, the father's age would have been
f-4 and it also would have been 9(x-4).
This will give you the equation: f-4 = 9(x-4)
- Substitute 'f' for 5x as shown above to give you your final equation: 5x-4 = 9(x-4)
4x = 32
x= 8
8 is the son's age now because it is equal to x and 5x8 is equal to the father's age now.
FINAL ANSWER: The father's current age is 40.